The displacement of a particle which moves along the s-axis is given by s = (t-3)exp(-0.6t)‚ where s is in meters and t is in seconds. Plot the displacement, velocity, and acceleration versus time for the first 20 seconds of motion. Determine the time at which the acceleration is zero?
Answer #1

Soft-3) e-0.67 At tzo -) S =-3 At t= f = S =-1.0976 At t22) Sa. 2-) So -0.3 At t = 3 ) S = Att zsz) Sz0.1 At tzlo S zo.0174 A-0.3 6.1 Sulst .o 6t e 0.60 e (4-3) (-0.63e obt Vee-oot -0.6t 0.6t Now. Va vist S = (t -3) Preds. dt [1 + (1 - 3) (-0.6) 0.61+ -0.61 Now v2.8 -0.6t 6476 a- dv 0.61 0.6e 0.61728-067 20-060 dt -0.6 e [-06-188+0.34+) Q: 62.28.70.36t)e-06 At t=0>) V-2:28